OpenStudy (anonymous):

x=3,-2,1 write a polynomial function in standard form with the given zeros

OpenStudy (anonymous):

But this is almost like the previous question, just that the numbers have changed! :(

OpenStudy (campbell_st):

well is x = 3 is a zero then (x - 3) is a binomial factor of the polynomial. just find the other binomial factors... then P(x) = a(x -3)(binomial factor 2)(binomial factor 3) a is a constant... based on your question I'd use a = 1

OpenStudy (anonymous):

i know but i got to (x^2-x-6)(x-1) n idk wat to do now

OpenStudy (campbell_st):

well its simply \[P(x) = x(x^2 -x -6) -1(x^2 -x -6)\] distribute and collect like terms to simplify

OpenStudy (anonymous):

how do u do it

OpenStudy (campbell_st):

well the 1st part is \[x^2 \times x -x \times x - 6 \times x\] and a similar process for the 2nd part.

OpenStudy (anonymous):

huh can u put the numbers tho

OpenStudy (campbell_st):

well what do you get \[x^2 \times x = \]

OpenStudy (anonymous):

i got (x-3)(x+2)(x-1) (x^2+2x-3x-6)(x-1) (x^2-1x-6)(x-1) wat do i do after that

OpenStudy (campbell_st):

yep thats correct.. and the polynomial is degree 3 so you need to distribute \[P(x) = (x^2 -x -6) \times(x -1)\] which can be done by splitting the binomial and multiplying \[P(x) = x \times (x^2 -x -6) -1\times(x^2 - x -6)\] you just need to simplify this.

OpenStudy (campbell_st):

once its multiplied out... collect the like terms.